Solve.
Find the length of the diagonal of a rectangle shaped book that has a length of 6 inches and width of 5 inches.

Answer by Jstrasner(112)

Hey,
So for this one, we need to use the pythagorean theorem:
a^2 + b^2 = c^2
Therefore:
6^2 + 5^2 = c^2
36 + 25 = 61
61 = c^2
c = 7.81
Therefore:
The length of the diagonal is about 8 inches
I hope this helps!